4.6/2. A thief steals an ATM card and must randomly guess the correct five​-digit pin code from a 5​-key keypad. Repetition of digits is allowed. What is the probability of a correct guess on the first​ try?

The number of possible codes is___.

​(Type an integer or fraction. Simplify your​ answer.)

the probability that correct code is given on the first try is?

​(Type an integer or fraction. Simplify your​ answer.)

3. If you know the names of the remaining seven students in the spelling​ bee, what is the probability of randomly selecting an order and getting the order that is used in the spelling​ bee?

​P(selecting the correct spelling bee ​order)=____

​(Type an integer or a simplified​ fraction.)

4. A fan of country music plans to make a custom CD with 11 of her 25 favorite songs. How many different combinations of 11 songs are​ possible? Is it practical to make a different CD for each possible​combination?

How many different combinations of 11 songs are​ possible?_____
Answer #1

2.

Since we have an 5-key keypad, and pin code is five digits so number of possible codes is 5^5 = 3125

Probability of guessing right on the first try is 1/3125= 0.00032

----------------------------------------------------------------------------------------------------------------------------

3.

We have 7 choices for the first student x 6 choices for the 2nd student x 5 choices for the 3rd student and so on.

So, the total number of possible arrangement is 7! = 5040

The probability of picking the correct order randomly would be 1 / 5040 = 0.0001984127

---------------------------------------------------------------------------------------------------------------------------